OpenStudy (anonymous):

simplify the expression. -2(2y-7x)-7=

OpenStudy (anonymous):

\[-4y+14x-7\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

that's the simplified form of this equation

OpenStudy (anonymous):

it says i can not factor

OpenStudy (anonymous):

What you do is distribute the -2 into the parentheses

OpenStudy (anonymous):

Distributive property : a(b + c) = ab + ac

OpenStudy (anonymous):

what do you mean you can't factor?

OpenStudy (anonymous):

So therefore -2(2y-7x)= -4y+14x

OpenStudy (anonymous):

Nice Assist Sean :)

OpenStudy (anonymous):

I try :P

OpenStudy (anonymous):

it says that not correct.

OpenStudy (anonymous):

what's the correct answer maybe i can figure out how they did it

OpenStudy (anonymous):

Trying to figure that out now.

OpenStudy (anonymous):

how do you know that this isn't correct, If you don't know the correct answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

because it pop up saying it......

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

is it \[y=\frac{ -7 }{ 2 }x+\frac{ 7 }{ 4 }\]

OpenStudy (anonymous):

i mean \[y=\frac{ 7 }{ 2 }x-\frac{ 7 }{ 4 }\]